# Singularities of symmetric powers and irrationality of motivic zeta functions

### v1

@@ 3. Irrationality theorem
In this section, we prove our main result on the irrationality of motivic zeta functions (Theorem 3.3), generalizing the method used in [LL2]. We first recall the construction of motivic measures $\mu_{k}$ which are used in the proof.
@@
Theorem 3.3.
Assume that at least one of the following conditions hold:
  • (a)
  • (b)
    There exists $i>0$ such that $H^{0}(X,\Omega_{X}^{2i})\neq 0$.
Then $Z(X,t)$ is not pointwise rational  [^6] For part (a), it is in fact enough to assume that $X$ is over an arbitrary field $K$ of characteristic zero. Indeed, using that $\operatorname{Sym}^{n}(X_{\overline{K}})=\operatorname{Sym}^{n}(X)_{\overline{K}}$, one reduces the proof to the case $K=\overline{K}$. One then checks that the proofs of all the lemmas which are used for part (a) work over such $K$..
We will need several lemmas.
@@
for all $i\geqslant 1$.
@@
By [AA, Proposition 1], $\operatorname{Sym}^{m}(X)$ has canonical singularities for all $m$. In particular, they have at most log terminal singularities and the statement follows from Lemma 3.7.
@@ 4. Some applications
For an irreducible variety $X$ over $K$ we denote by $\langle X\rangle$ its class in $\mathbb{Z}[\operatorname{SB}/K]$. Note that if $X$ is stably birational to $Y$, then $\operatorname{Sym}^{n}(X)$ is stably birational to $\operatorname{Sym}^{n}(Y)$ for all n>0n>0777Indeed, it is enough to show that $\operatorname{Sym}^{n}(X\times\mathbb{A}^{1})$ is birational to $\operatorname{Sym}^{n}(X)\times\mathbb{A}^{n}$. This follows from the fact the the projection $\rho_{1}:\operatorname{Sym}^{n}(X\times\mathbb{A}^{1})\to\operatorname{Sym}^{n}(X)$ is a locally trivial fibration over $X_{(1^{n})}\subset\operatorname{Sym}^{d}(X)$ (see notation 2.14) with fiber $\mathbb{A}^{n}$.. Therefore, we can define a formal power series:
@@
for every smooth connected projective variety $X$ and every $n>0$. In particular, we see that for such $X$ the image of $Z(X,t)$ in $(K_{0}(\mathcal{V}_{K})/(\mathbb{L}))[[t]]$ can be identified with $Z_{\operatorname{SB}}(X,t)$ using the isomorphism $\Phi$.
@@
The following is a generalization of Theorem 19 from [Litt2] on stable birationality of symmetric powers.
@@
Let us take $k$ as in the proof of Theorem 3.3, and $g_{s}\vcentcolon=\mu_{k}(\widetilde{\operatorname{Sym}}^{s}(X))$. Then, if the assumptions of the proposition hold, the polynomials $g_{m}$ and $g_{l}$ have different degrees (which was shown in the course of the proof of Theorem 3.3). In particular, $g_{m}\neq g_{l}$, and so $\widetilde{\operatorname{Sym}}^{m}(X)]\neq[\widetilde{\operatorname{Sym}}^{l}(X)]$ in $K_{0}(\mathcal{V}_{K})/(\mathbb{L})$, which is a contradiction. ∎
@@
By Proposition 3.6, for every smooth connected complex projective variety $X$ we have $H^{0}(Y_{m},\Omega^{i}_{Y_{m}})=H^{0}(X^{m},\Omega^{i}_{X^{m}})^{S_{m}}$ for all $m,i\geqslant 1$ [^8] Technically, in Proposition 3.6 in we assumed that $\dim(X)>1$, but it is straightforward to see that the equality also holds if $\dim(X)\leqslant 1$.. Moreover, $\mu_{1}(\lambda^{m}(X))=\mu_{1}(Y_{m})$ by Corollary 2.17, so we need to prove that

### v2

@@ 3. Irrationality theorem
In this section, we prove Theorem 1.2, building upon the method used in [LL2]. We first recall the construction of motivic measures $\mu_{k}$ which are used in the proof.
@@
To prove Theorem 1.2, we will need several lemmas, most of which already appear in [LL2] for the case when $X$ is a surface.
@@
for all $i\geqslant 1$ [^5] In fact, a similar argument shows that $H^{0}(\widetilde{Y},\Omega^{i}_{\widetilde{Y}})=H^{0}(Y,\Omega_{Y}^{[i]})=H^{0}(Z,\Omega_{Z}^{i})^{G}$, where $\Omega_{Y}^{[i]}$ is a sheaf of reflexive $i$-forms..
@@
By [AA, Proposition 1], $\operatorname{Sym}^{m}(X)$ has canonical singularities for all $m$, so the statement follows from Lemma 3.7.
@@
For part (a) of Theorem 1.2, it is in fact enough to assume that $X$ is over an arbitrary field $K$ of characteristic zero. Indeed, using that $\operatorname{Sym}^{n}(X_{\overline{K}})=\operatorname{Sym}^{n}(X)_{\overline{K}}$, one reduces the proof to the case $K=\overline{K}$. One then checks that the proofs of all the lemmas which are used for part (a) work over such $K$.
Remark 3.11.
@@ 4. Some applications
For an irreducible variety $X$ over $K$ we denote by $\langle X\rangle$ its class in $\mathbb{Z}[\operatorname{SB}/K]$. Note that if $X$ is stably birational to $Y$, then $\operatorname{Sym}^{n}(X)$ is stably birational to $\operatorname{Sym}^{n}(Y)$ for all $n>0$. Indeed, it is enough to show that $\operatorname{Sym}^{n}(X\times\mathbb{A}^{1})$ is birational to $\operatorname{Sym}^{n}(X)\times\mathbb{A}^{n}$, which follows from the fact the the projection $\rho_{1}:\operatorname{Sym}^{n}(X\times\mathbb{A}^{1})\to\operatorname{Sym}^{n}(X)$ is a locally trivial fibration over $X_{(1^{n})}\subset\operatorname{Sym}^{d}(X)$ (see Notation 2.12) with fiber $\mathbb{A}^{n}$ by [Göt, Lemma 4.4].
Hence, we can define a formal power series:
@@
for every smooth connected projective variety $X$ and every $n>0$. In particular, for such $X$ the image of $Z(X,t)$ in $(K_{0}(\mathcal{V}_{K})/(\mathbb{L}))[[t]]$ can be identified with $Z_{\operatorname{SB}}(X,t)$ using the isomorphism $\Phi$.
@@
The following is a generalization of Theorem 19 from [Litt] on stable birationality of symmetric powers.
@@
Let us take $k$ as in the proof of Theorem 1.2, and let $g_{s}\vcentcolon=\mu_{k}(\widetilde{\operatorname{Sym}}^{s}(X))$. Then, if the assumptions of the proposition hold, the polynomials $g_{m}$ and $g_{l}$ have different degrees (which was shown in the course of the proof of Theorem 1.2). In particular, $g_{m}\neq g_{l}$, and so $\widetilde{\operatorname{Sym}}^{m}(X)]\neq[\widetilde{\operatorname{Sym}}^{l}(X)]$ in $K_{0}(\mathcal{V}_{K})/(\mathbb{L})$, which is a contradiction. ∎
@@
By Proposition 3.6, for every smooth connected complex projective variety $X$ we have $H^{0}(Y_{m},\Omega^{i}_{Y_{m}})=H^{0}(X^{m},\Omega^{i}_{X^{m}})^{S_{m}}$ for all $m,i\geqslant 1$ [^6] Technically, in Proposition 3.6 we assumed that $\dim(X)>1$, but it is straightforward to see that the equality also holds if $\dim(X)\leqslant 1$.. Moreover, $\mu_{1}(\lambda^{m}(X))=\mu_{1}(Y_{m})$ by Corollary 2.15, so we need to prove that